Please set up Loan Amortization Schedules for the following question: You borrow $2000 at the rate of 12%, compounding monthly. You need to make monthly payment for three years.
Answer #1
Loan amortization Schedule:
Month ended Beginning Loan Interest Monthly Payment Reduction of loan Ending Loan
a b=a*12%*1/12 c d=c-b e=a-d
1 $ 2,000.00 $       20.00 $       66.43 $       46.43 $ 1,953.57
2 $ 1,953.57 $       19.54 $       66.43 $       46.89 $ 1,906.68
3 $ 1,906.68 $       19.07 $       66.43 $       47.36 $ 1,859.32
4 $ 1,859.32 $       18.59 $       66.43 $       47.84 $ 1,811.48
5 $ 1,811.48 $       18.11 $       66.43 $       48.31 $ 1,763.17
6 $ 1,763.17 $       17.63 $       66.43 $       48.80 $ 1,714.37
7 $ 1,714.37 $       17.14 $       66.43 $       49.28 $ 1,665.09
8 $ 1,665.09 $       16.65 $       66.43 $       49.78 $ 1,615.31
9 $ 1,615.31 $       16.15 $       66.43 $       50.28 $ 1,565.03
10 $ 1,565.03 $       15.65 $       66.43 $       50.78 $ 1,514.25
11 $ 1,514.25 $       15.14 $       66.43 $       51.29 $ 1,462.97
12 $ 1,462.97 $       14.63 $       66.43 $       51.80 $ 1,411.17
13 $ 1,411.17 $       14.11 $       66.43 $       52.32 $ 1,358.85
14 $ 1,358.85 $       13.59 $       66.43 $       52.84 $ 1,306.01
15 $ 1,306.01 $       13.06 $       66.43 $       53.37 $ 1,252.64
16 $ 1,252.64 $       12.53 $       66.43 $       53.90 $ 1,198.74
17 $ 1,198.74 $       11.99 $       66.43 $       54.44 $ 1,144.30
18 $ 1,144.30 $       11.44 $       66.43 $       54.99 $ 1,089.31
19 $ 1,089.31 $       10.89 $       66.43 $       55.54 $ 1,033.78
20 $ 1,033.78 $       10.34 $       66.43 $       56.09 $     977.69
21 $     977.69 $          9.78 $       66.43 $       56.65 $     921.04
22 $     921.04 $          9.21 $       66.43 $       57.22 $     863.82
23 $     863.82 $          8.64 $       66.43 $       57.79 $     806.03
24 $     806.03 $          8.06 $       66.43 $       58.37 $     747.66
25 $     747.66 $          7.48 $       66.43 $       58.95 $     688.71
26 $     688.71 $          6.89 $       66.43 $       59.54 $     629.17
27 $     629.17 $          6.29 $       66.43 $       60.14 $     569.03
28 $     569.03 $          5.69 $       66.43 $       60.74 $     508.29
29 $     508.29 $          5.08 $       66.43 $       61.35 $     446.94
30 $     446.94 $          4.47 $       66.43 $       61.96 $     384.99
31 $     384.99 $          3.85 $       66.43 $       62.58 $     322.41
32 $     322.41 $          3.22 $       66.43 $       63.20 $     259.20
33 $     259.20 $          2.59 $       66.43 $       63.84 $     195.37
34 $     195.37 $          1.95 $       66.43 $       64.47 $     130.89
35 $     130.89 $          1.31 $       66.43 $       65.12 $       65.77
36 $       65.77 $          0.66 $       66.43 $       65.77 $         0.00
Total $     391.43 $ 2,391.43 $ 2,782.86
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.01)^-36)/0.01 i = 1%
= 30.107505 n = 36
Monthly payment = Loan amount / Present value of annuity of 1
= $ 2,000.00 / 30.107505
= $       66.43